40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1850     Entry time: Thu Aug 3 19:59:10 2017
Author: awade 
Type: Misc 
Category: FSS 
Subject: Positive HV supply to FSS showing no current 

Despite the Craig's OLG TF looking ok for the north/south common path TF, the lock is breaking out into ringing with very small disturbances on the table.  Its not clear whats going on here. I through maybe either the PZT or EOM actuators were railing somewhere. The rms noise of the north error signal is typically around 18 mV on the north error signal monitor and 130 mV on the south: something is definitely not right there. this may be the previous problem that we had with south ringing.

I was looking around at the power supplies (just to check) and noticed that the positive +200 Kepco is registering no current on the front panel.  This could be the dial but it could also indicate that maybe we're missing on rail of the HV. This would be a problem for the EOM actuation stage of the FSS loop. 

Not sure how I'm going to test. I might open up the field boxes and stick a multi meter in there to see if +ve HV is actually getting there.

Attachment 1: 2017-08-03_19.47.36.jpg  2.285 MB  Uploaded Thu Aug 10 21:30:03 2017  | Hide | Hide all
2017-08-03_19.47.36.jpg
ELOG V3.1.3-